Oxygen gas (O2) is stored in steel cylinders. Assume we have 95.4 mol O2 stored in a 28-L cylinder at 25 oC. Use the van der Waals equation to calculate the pressure (in bar). [for O2(g), a = 137.8 and b = 0.0318]
Answer #1

n = 95.4 mole

V = 28L

T = 25 + 273 = 298K

R = 8.314*10^-2 L-bar/mole-K

a = 1.378bar-L^2/mole^2

b = 0.0318L/mole

(P + an^2/v^2)(V-nb) = nRT

(P + 1.378(95.4/28)^2)(28-95.4*0.0318) = 95.4*8.314*10^-2 *298

(P+ 15.99668)(24.96628) = 95.4*8.314*10^-2 *298

(P+ 15.99668)     = 95.4*8.314*10^-2 *298/(24.96628)

(P+ 15.99668)   = 94.67

P = 78.67bar